40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed Jul 14 14:39:09 2010, Dmass, Laser, PSL, H1 NPRO Measured NPROWaist.pdf
    Reply  Thu Jul 15 15:41:40 2010, Dmass, Laser, PSL, H1 NPRO Measured 
       Reply  Fri Jul 16 18:54:57 2010, Dmass, Laser, PSL, H1 NPRO Measured H1ModeMSoln.pdf
          Reply  Sat Jul 17 01:41:20 2010, Koji, Laser, PSL, H1 NPRO Measured 
          Reply  Wed Jul 21 20:13:06 2010, Dmass, Laser, PSL, H1 NPRO Measured H1NPRO_AM.png
             Reply  Thu Jul 22 16:19:48 2010, Koji, Laser, PSL, H1 NPRO Measured 
             Reply  Fri Jul 23 00:45:46 2010, Dmass, Laser, PSL, H1 NPRO Measured 
                Reply  Tue Jul 27 00:10:20 2010, Dmass, Laser, PSL, PMC Calibrated PMCCtrl.pdf
             Reply  Tue Jul 27 19:04:35 2010, Dmass, Laser, PSL, H1 NPRO Measured NproPZTFMBode.pdf
                Reply  Wed Jul 28 19:51:10 2010, Dmass, Laser, PSL, H1 NPRO Measured IMG_0241.png
                   Reply  Fri Jul 30 16:48:11 2010, Dmass, Laser, PSL, H1 NPRO Measured 
                      Reply  Sun Aug 1 18:37:07 2010, Dmass, Laser, PSL, H1 NPRO Measured 
                   Reply  Mon Aug 2 00:46:13 2010, Dmass, Laser, PSL, H1 NPRO Measured Picture_2.png
    Reply  Mon Aug 2 22:23:36 2010, Dmass, Laser, PSL, H1 NPRO Measured H1MM2PMC.pdf
    Reply  Tue Aug 3 01:31:15 2010, Dmass, Laser, PSL, H1 NPRO Measured H1NPROWaist2.pdf
       Reply  Tue Aug 3 13:47:55 2010, Dmass, Laser, PSL, H1 NPRO Measured Npro2PMCmm.pdf
       Reply  Tue Aug 3 17:27:20 2010, Dmass, Laser, PSL, H1 NPRO Measured 
Message ID: 871     Entry time: Tue Jul 27 00:10:20 2010     In reply to: 870
Author: Dmass 
Type: Laser 
Category: PSL 
Subject: PMC Calibrated 

I could sweep a couple FSR using a function generator into the PMC. A few numbers:

  • PMC FSR - 713 MHz (From design doc)
  • Voltage at drive box for 1 FSR = 3.5 +/- 0.3 V
    • Error bars generated via nonlinearity: (I ran through three FSR on one sweep, with one of the FSR around 0V input, the two FSR were 3.2 and 3.8 Volts, repsectively).
  •  
  • Voltage at PZT with 0V into driver = 142 V
  • 10 Vpp triangle wave at drive box = 184 Vpp at PZT drive (measured using a 10 Vpp triangle wave out of a function generator, watching the PZT voltage with a Fluke)
  • Calibration at input to drive box - 202 +/- 20MHz / Volt
  • Calibration at PZT - 11.1 MHz +/- 1 MHz/ Volt

Spectra shall be calibrated henceforth, or at the very least, calibratable.

Added: Calibrated PMC control signal in Hz/rtHz- As this should be totally coherent with err, I should just be able to divide by the open loop transfer function between ERR and CTRL...(I guess my Hz -> Volts bit of the OLTF on the error signal needs to be measured to get this no matter what...so I would end up with a Volts ->Hz for my error signal anyways...)

  • Foton isn't letting me add filters for some reason. It didn't earlier today, and Aidan had to change permissions (via chmod) to some process (Which?), then I could. I left and came back, and I could no longer write to Foton again.
  • Note that I am feeding back on the temperature of the NPRO.
  • WHOOPS: For correct spectrum I need to apply 2 poles @ 1Hz, 2 zeros at 10 Hz, 1 pole at 72 Hz, and a unity DC gain...
Attachment 1: PMCCtrl.pdf  47 kB  Uploaded Tue Jul 27 02:09:25 2010  | Hide | Hide all
PMCCtrl.pdf
ELOG V3.1.3-